[Koji, Nic, Manasa]
Update from last night.
Koji and I realigned the green optics on the PSL to start working on the ALS.
We set on a beat note search. We couldn't find the beat note between any of the arm green transmission and the PSL green. All we could see was the beat between the X arm and the Y arm green leakage.
Since we had the beatnote between the 2 green transmission beams, we decided to scan the PSl temperature. We scanned the SLOW actuator adjust of PSL; but couldn't locate any beat note. The search will continue again today.